OpenStudy (kirbykirby):

How do you solve 9 - 16x² > 0?

OpenStudy (anonymous):

you can see that this is in the form of a^2-b^2=(a-b)(a+b) that is in this case 3^2-(4x)^2=(3-4x)(3+4x)>0 when will this be greater than 0?

OpenStudy (anonymous):

when both are either positive or negative, that is

OpenStudy (anonymous):

yes! So what you need to do is solve 3-4x>0 and 3+4x>0 plus 0>3-4x and 0>3+45

OpenStudy (anonymous):

-16x^2>9 16x^2<-9 divide both sides by 16 x<-9/16

OpenStudy (anonymous):

?

OpenStudy (anonymous):

no not right

OpenStudy (anonymous):

if you graph \[y=9-16x^2\] you have a parabola that is facing down. it is zero at \[-\frac{3}{4}\] and \[\frac{3}{4}\]

OpenStudy (anonymous):

it will be positive (above the x - axis) between the zeros and negative otherwise

OpenStudy (anonymous):

16 x^2<9 (x-3\/4) (x+3\/4)>0 -(4 x-3) (4 x+3)>0 -3/4<x<3/4

OpenStudy (anonymous):

you want to know where it is positive, so answer is ... what godefroid wrote : \[(-\frac{3}{4},\frac{3}{4})\]

OpenStudy (kirbykirby):

Thank you so much :)!